# OT Score: An OT based Confidence Score for Unsupervised Domain Adaptation

### v1

@@ 4.2 Post-check and OT Score
In practice, this $g$ value gap can be used as a post-check tool once target pseudo labels have been assigned by a DA algorithm. We can compute the gap $\inf_{x\in{\nu}_{1}}g(x)-\sup_{x\in{\nu}_{2}}g(x)$ based on pseudo-labeled partition of the target distribution ${\nu}_{1}$ and ${\nu}_{2}$. In addition to global assessment, the individual $g(x)$ values can also serve as confidence indicators. Specifically, for target samples pseudo-labeled as class $\nu_{2}$, larger $g(x)$ values indicate higher classification confidence; conversely, for samples labeled as class $\nu_{1}$, smaller $g(x)$ values indicate higher confidence.
@@
(1) Improved computational efficiency: Semi-discrete OT offers better computational complexity compared to fully discrete OT.
(2) Efficient incremental optimization: When a subset of data points is added or removed, the optimization of semi-discrete OT can be efficiently continued using stochastic gradient descent (SGD), without requiring a full recomputation from scratch.
@@
The following corollary might be helpful in some computation scenarios.
@@ 5 OT score Computation
In this section, we present the algorithm used to compute the OT score and extend the definition to multiclass setting. Specifically, we model the source distribution in the feature space as a discrete measure and treat the target data as samples drawn from a continuous measure. Our approach is developed under the few-shot Source-Free UDA setting, where only very limited source information of feature representations is available. In the experimental section, we demonstrate that the OT score achieves strong performance even when using only the class-wise mean feature representations from the source domain. This not only reduces the computational complexity but also addresses privacy and security concerns associated with access to full source data.
Definition 1.
where $\tilde{d}(\cdot,\cdot)$ requires computing the semi-discrete OT. And we define the OT score on target samples as
$g(x):=\min_{j}g_{j}(x).$
We summarize our OT score computation in Algorithm 1.
@@
1:Input: Source feature representations $\mathbf{Z}^{\mathcal{S}}$, corresponding labels $\bm{y}^{\mathcal{S}}$, source sample weights $\bm{a}$, target features $\mathbf{Z}^{\mathcal{T}}$ and corresponding predicted labels (or pseudo labels) $\hat{\bm{y}}^{\mathcal{T}}$, entropic regularization parameter $\varepsilon$, learning rate $\gamma$.
@@
Remark 4.
Inspired by the OT score, we propose SDOT-DA, a simple yet effective algorithm for UDA. It assigns labels to targets based on the semi-discrete OT map from target features to source features, using the labels of their transport destinations. Further details are provided in Appendix C.
6 Experiemnts
In this section, we present experiments to evaluate the effectiveness of our algorithms. Additional details and results are provided in Appendix D.
6.1 Synthetic Data
In this section, we use synthetic data to validate and illustrate our theoretical findings. Specifically, we consider a 2D scenario where data points are sampled from circular regions. The source domain consists of class-separated samples drawn from disjoint circles, whereas the target domain includes clusters with partial overlap. The distribution of the generated data is visualized in Figure LABEL:fig:overlap_samples. We compute the max-min values as described in Theorem 6 and present the results in Figure LABEL:fig:overlap_true. As shown in Figure LABEL:fig:overlap_predict, the SDOTDA algorithm produces label predictions in which many samples within the overlapping regions are misclassified. However, after removing low-confidence predictions, the remaining samples are almost entirely classified correctly, as illustrated in Figure LABEL:fig:overlap_pred_quantil_add. Notably, the separation between the two clusters becomes significantly more obvious after this filtering step. Additional binary digit classification results under added noise conditions are provided in Appendix D.
6.2 Real World Datasets
@@
For evaluation, we adopt the Area Under the Risk-Coverage Curve (AURC) proposed by [7] [5] and subsequently employed in [15]. Specifically, after obtaining the high-confidence subset $X_{h}^{\mathcal{T}}:=\left\{x_{i}^{\mathcal{T}}\mid s\left(x_{i}^{\mathcal{T}},\hat{y}_{i}^{\mathcal{T}}\right)>h\right\}$, where $h$ is a predefined confidence threshold, the risk is computed as the average empirical loss over $X_{h}^{\mathcal{T}}$, and the coverage corresponds to $\left|X_{h}^{\mathcal{T}}\right|/\left|X^{\mathcal{T}}\right|$. A lower AURC value indicates higher confidence reliability, as it implies a lower prediction risk at a given coverage level. Notably, when the $0/1$ loss is applied, a high AURC reflects a high error rate among pseudo-labels, thus indicating poor correctness and calibration of the confidence scores.
To further assess the robustness of the proposed OT score under varying pseudo-label quality, we consider two evaluation scenarios: (1) pseudo labels obtained using a Gaussian Mixture Model (GMM); and (2) pseudo labels generated by the DSAN algorithm [32] with 20 epochs of training.
Under setting (1), Maxprob and Ent uses labels assigned by the pretrained source classifier while Cossim, JMDS, OT score receive pseudo labels from a Gaussian Mixture Model (GMM). We also considered to assign pseudo labels using the SDOT-DA algorithm. However, empirical results showed that GMM-based labels yield better performance. Therefore, for the comparisons presented in Table 2, we adopt GMM-generated pseudo labels for computing the OT score. Under setting (2), only Cossim and OT score are capable of incorporating externally generated high-quality pseudo labels. We continue under the setting where only the mean feature representation of each source class is available. Table 1 shows the significant benefits of leveraging high quality pseudo labels. The OT score achieved the lowest AURC value in most adaptation tasks across the considered scenarios.
Table 2: Evaluation of confidence scores based on AURC (GMM).
Dataset Task Maxprob Ent Cossim JMDS OT Score
Digits M $\rightarrow$ U 0.0055 0.0049 0.0051 0.0053 0.0033
U $\rightarrow$ M 0.0224 0.0228 0.0204 0.0355 0.0214
Avg. 0.0140 0.0139 0.0128 0.0204 0.0124
Office-Home Ar $\rightarrow$ Cl 0.2823 0.2860 0.2871 0.2825 0.2777
Ar $\rightarrow$ Pr 0.1320 0.1353 0.1219 0.1328 0.1313
Ar $\rightarrow$ Rw 0.0782 0.0835 0.1073 0.0820 0.0810
Cl $\rightarrow$ Ar 0.2421 0.2450 0.2503 0.2971 0.2218
Cl $\rightarrow$ Pr 0.1489 0.1517 0.1336 0.1708 0.1354
Cl $\rightarrow$ Rw 0.1366 0.1384 0.1412 0.1595 0.1241
Pr $\rightarrow$ Ar 0.2355 0.2453 0.2580 0.2715 0.2425
Pr $\rightarrow$ Cl 0.3417 0.3419 0.3180 0.3390 0.3257
Pr $\rightarrow$ Rw 0.0836 0.0875 0.1007 0.0952 0.0781
Rw $\rightarrow$ Ar 0.1439 0.1487 0.1785 0.1750 0.1391
Rw $\rightarrow$ Cl 0.2826 0.2812 0.2762 0.2760 0.2613
Rw $\rightarrow$ Pr 0.0693 0.0713 0.0725 0.0890 0.0637
Avg. 0.1814 0.1846 0.1871 0.1971 0.1735
VisDA-2017 T $\rightarrow$ V 0.3327 0.3312 0.3494 0.3090 0.2916
@@ 7 Conclusion and Future Work
We investigate theoretical guarantees about allowed distribution shifts in order to have a label-preserving OT. Using semi-discrete OT, we derive the OT score which considers the decision boundary induced by the OT alignment. The definition of OT score can be easily extended to other cost functions other than the standard Euclidean norm. Additionally, confidence scores are helpful for sample reweighting during training process. Since optimization of semi-discrete OT can be efficiently continued using stochastic gradient descent (SGD), it is also useful to propose a recomputation scheme for OT score during training.
Also, in our experiments, we only use the source mean class features for the discrete measure, it remains to explore how to increase the source information in the most efficient way. Intuitively, we want to replace the source empirical distribution $\hat{\mu}$ by $\hat{\mu}_{new}$ defined on a smaller finite set such that $T_{\nu}^{\hat{\mu}}\approx T_{\nu}^{\hat{\mu}_{new}}$. We leave this investigation for future work.
Appendix A Confidence Scores
@@
$\begin{gathered}\operatorname{Maxprob}\left(x_{i}^{\mathcal{T}}\right)=\max_{c}p_{\mathcal{S}}\left(x_{i}^{\mathcal{T}}\right)_{c},\\ \operatorname{Ent}\left(x_{i}^{\mathcal{T}}\right)=1+\frac{\Sigma_{c=1}^{K}p_{\mathcal{S}}\left(x_{i}^{\mathcal{T}}\right)_{c}\log p_{\mathcal{S}}\left(x_{i}^{\mathcal{T}}\right)_{c}}{\log K},\\ \operatorname{Cossim}\left(x_{i}^{\mathcal{T}}\right)=\frac{1}{2}\left(1+\frac{\left\langle x_{i}^{\mathcal{T}},C_{\hat{y}_{i}^{\mathcal{T}}}\right\rangle}{\left\|x_{i}^{\mathcal{T}}\right\|\left\|C_{\hat{y}_{i}^{\mathcal{T}}}\right\|}\right).\\ \end{gathered}$
JMDS score is computed by $\operatorname{JMDS}\left(x_{i}^{\mathcal{T}}\right)=\operatorname{LPG}\left(x_{i}^{\mathcal{T}}\right)\cdot\operatorname{MPPL}\left(x_{i}^{\mathcal{T}}\right)$. LPG is the Log-Probability Gap computed from log data-structure-wise probability $\log p_{\text{data }}\left(x_{i}^{\mathcal{T}}\right)$ using GMM on the target feature space. MPPL provides high scores for samples whose GMM pseudo-label is the same based on $p_{\mathcal{S}}\left(x_{i}^{\mathcal{T}}\right)$ and $p_{\text{data }}\left(x_{i}^{\mathcal{T}}\right)$. Details about JMDS score can be found in [15].
Appendix B Proofs
Proof of Theorem 4.

### v2

@@ 3.2 Semi-Discrete Setting
In practice, this $g$ value gap can be used as a post-check tool once target pseudo labels have been assigned by any algorithm. We can compute the gap $\inf_{x\in{\nu}_{1}}g(x)-\sup_{x\in{\nu}_{2}}g(x)$ based on pseudo-labeled partition of the target distribution ${\nu}_{1}$ and ${\nu}_{2}$. In addition to global assessment, the individual $g(x)$ values can also serve as confidence indicators. Specifically, for target samples pseudo-labeled as class $\nu_{2}$, larger $g(x)$ values indicate higher classification confidence; conversely, for samples labeled as class $\nu_{1}$, smaller $g(x)$ values indicate higher confidence.
@@
(1) Efficient incremental optimization: Semi-discrete OT can be updated incrementally with SGD instead of being solved from scratch. As target pseudo-labels evolve, we reuse the previous solution as initialization and perform a few mini-batch SGD updates to reflect the new assignments.
@@
The following corollary might be helpful in some computation scenarios: it enables computing the semi-discrete OT for each component separately, thereby reducing the dimension of the dual weights.
@@ 4 OT score Computation
1:Input: Source class-wise mean feature representations $\mathbf{Z}^{\mathcal{S}}$, corresponding labels $\bm{y}^{\mathcal{S}}$, source sample weights $\bm{a}$, target features $\mathbf{Z}^{\mathcal{T}}$ and corresponding predicted labels (or pseudo labels) $\hat{\bm{y}}^{\mathcal{T}}$, entropic regularization parameter $\varepsilon$, learning rate $\gamma$.
@@
In this section, we extend the definition of OT score to multiclass setting and present the algorithm used for computation. Specifically, we model the source distribution in the feature space as a discrete measure and treat the target data as samples drawn from a continuous measure.
Definition 1.
[…]
$g(x):=\min_{j}g_{j}(x).$
We summarize our OT score computation in Algorithm 1. We represent the source distribution by class-wise mean features. Accordingly, the definition of $g_{j}$ simplifies to $g_{j}(x)=\tilde{d}(x,\mu_{j})-\tilde{d}(x,\mu_{i})$, where $\mu_{i}$ and $\mu_{j}$ are the mean features of classes $i$ and $j$, respectively. Under this setting, we show that classification accuracy increases as samples with low OT scores are filtered out.
Theorem 8.
Let $\nu_{1},\nu_{2}$ be the continuous probability measures with means $m_{1}$ and $m_{2}$, respectively and $\hat{\mu}_{i}$ consists of singletons $y_{i}$. Denote $\nu:=\frac{1}{2}\nu_{1}+\frac{1}{2}\nu_{2}$ and $\hat{\mu}:=\frac{1}{2}\hat{\mu}_{1}+\frac{1}{2}\hat{\mu}_{2}$. Suppose $\nu_{i}(|X_{i}-m_{i}|\geq t)\leq 2\exp\left(-\frac{t^{2}}{2\sigma^{2}}\right)$ and $\|m_{1}-y_{1}\|+\|m_{2}-y_{2}\|<\|m_{1}-y_{2}\|+\|m_{2}-y_{1}\|$, then $P\Big(T_{\nu}^{\hat{\mu}}(X_{i})\neq Y_{i}|g(X_{i})>g\Big)\leq 2\exp\left(-\frac{\min_{i=1,2}\operatorname{dist}(m_{i},\mathcal{S})^{2}}{2\sigma^{2}}\right)$, where
(1) $\mathcal{S}:=\Bigl\{x:\|x-y_{1}\|-(w^{*}+g)=\|x-y_{2}\|\Bigr\}$
(2) $d:=\|y_{2}-y_{1}\|,\qquad e:=\frac{y_{2}-y_{1}}{d},\qquad$ $m=\alpha e+u,\;u\perp e$ is the orthogal decomposition of $m$ and denote $\rho:=\|u\|$.
[…]
$\sqrt{t^{2}+r^{2}}\;=\;\sqrt{(t-d)^{2}+r^{2}}+(w^{*}+g),\qquad r\geq 0.$
5 Applications and Empirical Evaluation
Table 1: Evaluation of confidence scores based on AURC.
Dataset Task Maxprob Ent Cossim JMDS OT Score
Office‑Home Ar $\rightarrow$ Cl 0.3485 0.3592 0.3013 0.2885 0.2623
Ar $\rightarrow$ Pr 0.1697 0.1789 0.1297 0.1237 0.1208
Ar $\rightarrow$ Rw 0.1032 0.1133 0.0897 0.0797 0.0770
Cl $\rightarrow$ Ar 0.2686 0.2849 0.2045 0.2362 0.2020
Cl $\rightarrow$ Pr 0.1916 0.2027 0.1182 0.1483 0.1424
Cl $\rightarrow$ Rw 0.1703 0.1837 0.1180 0.1275 0.1179
Pr $\rightarrow$ Ar 0.2629 0.2753 0.1977 0.2123 0.2063
Pr $\rightarrow$ Cl 0.3910 0.4052 0.3189 0.3193 0.3249
Pr $\rightarrow$ Rw 0.0997 0.1085 0.0757 0.0786 0.0741
Rw $\rightarrow$ Ar 0.1516 0.1621 0.1315 0.1369 0.1167
Rw $\rightarrow$ Cl 0.3339 0.3463 0.2873 0.2664 0.2539
Rw $\rightarrow$ Pr 0.0731 0.0796 0.0639 0.0737 0.0557
Avg. 0.2137 0.2250 0.1697 0.1743 0.1628
VisDA‑2017 T $\rightarrow$ V 0.3071 0.3203 0.2780 0.2021 0.1704
ImageCLEF-DA C $\rightarrow$ I 0.0515 0.0570 0.0181 0.0325 0.0252
C $\rightarrow$ P 0.1902 0.1991 0.1579 0.1459 0.1143
I $\rightarrow$ C 0.0099 0.0131 0.0038 0.0055 0.0036
I $\rightarrow$ P 0.1198 0.1221 0.1280 0.1170 0.1000
P $\rightarrow$ C 0.0260 0.0303 0.0062 0.0216 0.0092
P $\rightarrow$ I 0.0347 0.0382 0.0177 0.0276 0.0186
Avg. 0.0720 0.0766 0.0553 0.0583 0.0452
In this section, we present: (i) an Area Under the Risk–Coverage Curve (AURC) evaluation across confidence scores (Section 5.1); (ii) an SFUDA application using the OT score for training-time reweighting to improve accuracy (Section 5.2); and (iii) a label-free model-selection analysis showing that the mean OT score on the target set correlates with final accuracy (Section 5.3). Additional details and results are provided in Appendix D.
5.1 AURC Comparisons
@@
For evaluation, we adopt the Area Under the Risk-Coverage Curve (AURC) proposed by Geifman et al. (2018); Ding et al. (2020) and subsequently employed in Lee et al. (2022). Specifically, after obtaining the high-confidence subset $X_{h}^{\mathcal{T}}:=\left\{x_{i}^{\mathcal{T}}\mid s\left(x_{i}^{\mathcal{T}},\hat{y}_{i}^{\mathcal{T}}\right)>h\right\}$, where $h$ is a predefined confidence threshold, the risk is computed as the average empirical loss over $X_{h}^{\mathcal{T}}$, and the coverage corresponds to $\left|X_{h}^{\mathcal{T}}\right|/\left|X^{\mathcal{T}}\right|$. A lower AURC value indicates higher confidence reliability, as it implies a lower prediction risk at a given coverage level. Notably, when the $0/1$ loss is applied, a high AURC reflects a high error rate among pseudo-labels, thus indicating poor correctness and calibration of the confidence scores.
Maxprob and Ent use labels assigned by the pretrained source classifier while Cossim, JMDS, OT score receive pseudo labels from a Gaussian Mixture Model (GMM), following the same setup of Lee et al. (2022).
To further assess the robustness of the proposed OT score under varying pseudo-label quality, we consider another case where the pseudo labels are generated by the DSAN algorithm (Zhu et al., 2020). Under this setting, only Cossim and OT score are capable of incorporating externally generated high-quality pseudo labels. Table 5 in Appendix D shows the significant benefits of leveraging high quality pseudo labels. The OT score achieves the lowest AURC value in most adaptation tasks across the considered scenarios.
5.2 OT Score Reweighting
We integrate the OT score into CoWA-JMDS (Lee et al., 2022) as a per-sample weight for pseudo-labeled target instances. For each target sample $x_{i}$, we set
$w_{i}\;=\;2\cdot\mathrm{OT}(x_{i})\,\cdot\mathrm{JMDS}(x_{i}),$
where $\mathrm{JMDS}(x_{i})$ is computed online from features during training, while $\mathrm{OT}(x_{i})$ is computed from features extracted by the pre-adaptation model, thereby decoupling the confidence signal from the evolving target representation. Relying solely on the same training-time features that are continually updated by pseudo-labels risks self-reinforcement (confirmation bias): incorrect pseudo-labels $\rightarrow$ representation drift $\rightarrow$ inflated “confidence” $\rightarrow$ further amplification. We mitigate this by computing the OT score from pre-adaptation features, which constrains the pseudo-label feedback loop and reduces confirmation bias. Here, the OT score is normalized to $[0,1]$; the prefactor $2$ offsets the dynamic-range compression induced by the product of two numbers in $[0,1]$.
This integration yields higher accuracy than the original CoWA-JMDS. We evaluate on Office-Home (Tables 2) and VisDA-2017 (Tables 3) in the SFUDA setting, reporting target-domain accuracy averaged over three seeds (see Appendix D).Training settings (backbone, optimizer, pseudo-labeling) follow Lee et al. (2022); the only change is the per-sample weight $w_{i}$.
[…]
BAIT (Yang et al., 2020) 57.4 77.5 82.4 68.0 77.2 75.1 67.1 55.5 81.9 73.9 59.5 84.2 71.6
SHOT (Liang et al., 2020) 57.1 78.1 81.5 68.0 78.2 78.1 67.4 54.9 82.2 73.3 58.8 84.3 71.8
NRC (Yang et al., 2021) 57.7 80.3 82.0 68.1 79.8 78.6 65.3 56.4 83.0 71.0 58.6 85.6 72.2
ELR (Yi et al., 2023) 58.4 78.7 81.5 69.2 79.5 79.3 66.3 58.0 82.6 73.4 59.8 85.1 72.6
CPD (Zhou et al., 2024) 59.1 79.0 82.4 68.5 79.7 79.5 67.9 57.9 82.8 73.8 61.2 84.6 73.0
CoWA (Lee et al., 2022) 56.9 78.4 81.0 69.1 80.0 79.9 67.7 57.2 82.4 72.8 60.5 84.5 72.5
OTScore 58.0 79.6 81.5 69.6 80.2 80.0 68.3 57.6 82.3 73.2 61.1 84.7 73.0
Table 2: Accuracy (%) on Office-Home (ResNet-50).
Method plane bcycl bus car horse knife mcycl person plant sktbrd train truck Avg
SFIT (Hou & Zheng, 2021) 94.3 79.0 84.9 63.6 92.6 92.0 88.4 79.1 92.2 79.8 87.6 43.0 81.4
SHOT (Liang et al., 2020) 94.3 88.5 80.1 57.3 93.1 94.9 80.7 80.3 91.5 89.1 86.3 58.2 82.9
NRC (Yang et al., 2021) 96.8 91.3 82.4 62.4 96.2 95.9 86.1 80.6 94.8 94.1 90.4 59.7 85.9
AdaCon (Chen et al., 2022) 97.0 84.7 84.0 77.3 96.7 93.8 91.9 84.8 94.3 93.1 94.1 49.7 86.8
ELR (Yi et al., 2023) 97.3 89.1 89.8 79.2 96.9 97.5 92.2 82.5 95.8 94.5 87.3 34.5 86.4
CPD (Zhou et al., 2024) 96.7 88.5 79.6 69.0 95.9 96.3 87.3 83.3 94.4 92.9 87.0 58.7 85.8
CoWA (Lee et al., 2022) 96.2 89.7 83.9 73.8 96.4 97.4 89.3 86.8 94.6 92.1 88.7 53.8 86.9
OTScore 95.6 89.0 82.8 78.3 96.3 98.0 91.2 86.8 95.5 94.7 89.9 55.7 87.8
Table 3: Accuracy (%) on VisDA-2017 (ResNet-101).
5.3 Model Comparison
[…]
Figure 2: Mean OT Score vs. accuracy on Office-Home. Lines connect targets sharing the same source. Points denote individual target domains.
Table 4: Accuracy (%) on USPS with different sources.
Source Mean Score Accuracy (%)
MNIST 4.02 94.7
FLIP-USPS 0.55 47.8
@@ 6 Conclusion and Future Work
We investigate theoretical guarantees about allowed distribution shifts in order to have a label-preserving OT. Using semi-discrete OT, we derive the OT score which considers the decision boundary induced by the OT alignment. The definition of OT score can be easily extended to other cost functions other than the standard Euclidean norm. Additionally, confidence scores are helpful for training-time sample reweighting and model comparison.
Currently, we address class imbalance in the OT-score computation by weighting the source class mean features with class proportions estimated from pseudo labels. However, when pseudo labels are unreliable, these estimates can be biased. Under the assumptions in Section E, we show that the OT objective is minimized when the source and target class proportions coincide (see Theorem 11). A natural next step is to model and propagate class-proportion uncertainty into the confidence score.
Appendix A Related Works
Theory about DA: Several theoretical works have investigated the learnability and generalization guarantees of domain adaptation (DA). Specifically, Ben-David & Urner (2012) analyzes the DA learnability problem and sample complexity under the standard VC-dimension framework, and identifies a setting in which no algorithm can successfully solve the DA problem. In a related direction, Redko et al. (2019) provides a theoretical analysis about the existence of a hypothesis that performs well across both source and target domains, and further establishes finite-sample approximation properties of the $\lambda$ term. Le et al. (2021) alleviates the label mismatching problem by searching for a transformation $T$ that satisfies the following conditions: (1) $T_{\#}\mu_{S}=\mu_{T}$, and (2) $T$ preserves the labels.
Confidence Scores: Uncertainty estimation and confidence score have been prevalently employed in machine learning to improve model robustness. In particular, ordinal ranking techniques have been commonly used for selective classification (Lakshminarayanan et al. (2017); Geifman & El-Yaniv (2017); Mandelbaum & Weinshall (2017); Nair et al. (2020)), where the goal is to prioritize or filter samples based on their confidence scores in order to exclude low-confidence samples during training. Karim et al. (2023) select reliable pseudo-labels by thresholding the maximum softmax probability of the teacher’s augmentation-averaged prediction. Litrico et al. (2023) reweight the classification loss by entropy, assigning higher weights to low-entropy (more confident) samples. Lee et al. (2022) propose the JMDS score to effectively identify low-confidence samples, thereby enhancing the reliability of the DA process. However, most existing confidence scores rely primarily on cluster-level information in the feature space, without explicitly modeling the geometric relationship between domains. In contrast, our proposed OT score take into account the geometry induced by the OT map, establishing a stronger connection between the source and target domains when computing confidence scores.
Appendix B Confidence Scores
@@
[…]
JMDS score is computed by $\operatorname{JMDS}\left(x_{i}^{\mathcal{T}}\right)=\operatorname{LPG}\left(x_{i}^{\mathcal{T}}\right)\cdot\operatorname{MPPL}\left(x_{i}^{\mathcal{T}}\right)$. LPG is the Log-Probability Gap computed from log data-structure-wise probability $\log p_{\text{data }}\left(x_{i}^{\mathcal{T}}\right)$ using GMM on the target feature space. MPPL provides high scores for samples whose GMM pseudo-label is the same based on $p_{\mathcal{S}}\left(x_{i}^{\mathcal{T}}\right)$ and $p_{\text{data }}\left(x_{i}^{\mathcal{T}}\right)$. Details about JMDS score can be found in Lee et al. (2022).
Appendix C Proofs
Proof of Theorem˜4.